OpenStudy (anonymous):

Using synthetic division, what is the quotient: (3x3 + 4x - 32) ÷ (x - 2)

OpenStudy (anonymous):

Solve for x: 3 x^2+6 x+16 = 0 Write the quadratic equation in standard form. Divide both sides by 3: x^2+2 x+16/3 = 0 Solve the quadratic equation by completing the square. Subtract 16/3 from both sides: x^2+2 x = -16/3 Take one half of the coefficient of x and square it, then add it to both sides. Add 1 to both sides: x^2+2 x+1 = -13/3 Factor the left hand side. Write the left hand side as a square: (x+1)^2 = -13/3 Eliminate the exponent on the left hand side. Take the square root of both sides: x+1 = i sqrt(13/3) or x+1 = -i sqrt(13/3) Look at the first equation: Solve for x. Subtract 1 from both sides: x = i sqrt(13/3)-1 or x+1 = -i sqrt(13/3) Look at the second equation: Solve for x. Subtract 1 from both sides: Answer: | | x = i sqrt(13/3)-1 or x = -i sqrt(13/3)-1